OpenStudy (anonymous):

Find the real numbers of x & y that make the equation true (5-2i) - 7= x-(3+yi)

OpenStudy (anonymous):

Is there a second equation?

OpenStudy (anonymous):

@_ambitiousiam you should isolating x and y...

OpenStudy (anonymous):

5 - 2i - 7 = x - 3 -yi 5 - 2i - 7 + 3 = x - yi 1-2i = x- yi so, x=1 and y=2